How many grams of CuO will be formed from the decomposition of 2.1 grams of Cu(OH)

2?


1
Expert's answer
2021-08-02T02:49:04-0400

Solution:

Balanced chemical equation:

Cu(OH)2 → CuO + H2O

According to the equation above: n(Cu(OH)2) = n(CuO)


Determine the moles of Cu(OH)2:

The molar mass of Cu(OH)2 is 97.561 g/mol.

Therefore,

[2.1 g Cu(OH)2] × [1 mol Cu(OH)2 / 97.561 g Cu(OH)2] = 0.0215 mol Cu(OH)2


Copper(II) hydroxide and copper(II) oxide are equivalent to 1:1.

Thus, n(CuO) = n(Cu(OH)2) = 0.0215 mol


Determine the mass of CuO:

The molar mass of CuO is 79.545 g/mol.

Therefore,

[0.0215 mol CuO] × [79.545 g CuO / 1 mol CuO] = 1.71 g CuO = 1.7 g CuO


Answer: 1.7 g of CuO will be formed.